There are 16 members on a city council. At a recent city council meeting, 9 of the council members voted in favor of a budget increase. How many possible groups of council members could have voted in favor?


Will the calculation involve permutations or combinations?

  • combinations
  • permutations

There are ?  possible groups of 9 city council members who voted in favor of the budget increase


1
Expert's answer
2021-09-21T02:57:26-0400

The order of the council members to vote in favor of a budget increase will not change the group of the members. That is to say, the order is not considered here

Therefore calculation involves combinations

There are "_{9}^{16}C" possible groups "\{_{r}^nC=\frac{n!}{(n-r)!*r!}\}"

"\implies _{9}^{16}C=\frac{16!}{(16-9)!*9!}=\frac{16!}{(7)!*9!}=11440"

There are 11440 possible groups of 9 city council members who voted in favor of the budget increase